During a 2-week period, the average number of accidents on this section of highway is 4. What is the probability of at most two accidents during a 2-week period?


1
Expert's answer
2021-02-24T06:32:19-0500

Let "X=" the number of  accidents on this section of highway during a

2-week period: "X \sim Po(\lambda)"


"P(X=x)=\dfrac{e^{-\lambda}\cdot\lambda^x}{x!}"

Given "\lambda=4"


"P(X\leq2)=P(X=0)+P(X=1)+P(X=2)"

"=\dfrac{e^{-4}\cdot4^0}{0!}+\dfrac{e^{-4}\cdot4^1}{1!}+\dfrac{e^{-4}\cdot4^2}{2!}"

"=13e^{-4}\approx0.238103"

The probability of at most two accidents during a 2-week period is 0.238103.